# EgoEvGesture: Gesture Recognition Based on Egocentric Event Camera

### v2

@@ Abstract
Egocentric gesture recognition is a pivotal technology for enhancing natural human-computer interaction, yet traditional RGB-based solutions suffer from motion blur and illumination variations in dynamic scenarios. While event cameras show distinct advantages in handling high dynamic range with ultra-low power consumption, existing RGB-based architectures face inherent limitations in processing asynchronous event streams due to their synchronous frame-based nature. Moreover, from an egocentric perspective, event cameras record data that includes events generated by both head movements and hand gestures, thereby increasing the complexity of gesture recognition. To address this, we propose a novel network architecture specifically designed for event data processing, incorporating (1) a lightweight CNN with asymmetric depthwise convolutions to reduce parameters while preserving spatiotemporal features, (2) a plug-and-play state-space model as context block that decouples head movement noise from gesture dynamics, and (3) a parameter-free Bins-Temporal Shift Module (BSTM) that shifts features along bins and temporal dimensions to fuse sparse events efficiently. We further establish the EgoEvGesture dataset, the first large-scale dataset for egocentric gesture recognition using event cameras. Experimental results demonstrate that our method achieves 62.7% accuracy tested on unseen subjects with only 7M parameters, 3.1% higher than state-of-the-art approaches. Notable misclassifications in freestyle motions stem from high inter-personal variability and unseen test patterns differing from training data. Moreover, our approach achieved a remarkable accuracy of 97.0% on the DVS128 Gesture, demonstrating the effectiveness and generalization capability of our method on public datasets. The dataset and models are made available at https://github.com/3190105222/EgoEv_Gesture.
@@ II-B Event-based Methods for Human Perception
TABLE I: Ultra-compact dataset comparison with key attributes
@@
Attributes: M: Modality (R=RGB, E=Event), V: Volume, C: Classes, 2H: Two-Hand, R: Resolution (pixels), Y: Year, W: Real-World, 1P: 1st-Person
@@ V-D Ablation Study
V-E Quantitatively Result on DVS128 Gesture
@@
Our leading recognition performance on the DVS128 Gesture dataset stems from three key architectural innovations: (1) A lightweight CNN with asymmetric depthwise convolutions that maintains spatiotemporal features while reducing parameters, crucial for sparse event-based data; (2) A plug-and-play state-space context block that focuses on gesture dynamics; (3) The parameter-free BSTM module that effectively fuses sparse spatiotemporal events. Achieving 97.0% accuracy (matching S-tllr’s reimplemented performance) validates our architecture’s generalization capability and robustness for event-based gesture recognition.

### v3

@@ Abstract
Egocentric gesture recognition is a pivotal technology for enhancing natural human-computer interaction, yet traditional RGB-based solutions suffer from motion blur and illumination variations in dynamic scenarios. While event cameras show distinct advantages in handling high dynamic range with ultra-low power consumption, existing RGB-based architectures face inherent limitations in processing asynchronous event streams due to their synchronous frame-based nature. Moreover, from an egocentric perspective, event cameras record data that includes events generated by both head movements and hand gestures, thereby increasing the complexity of gesture recognition. To address this, we propose a novel network architecture specifically designed for event data processing, incorporating (1) a lightweight CNN with asymmetric depthwise convolutions to reduce parameters while preserving spatiotemporal features, (2) a plug-and-play state-space model as context block that decouples head movement noise from gesture dynamics, and (3) a parameter-free Bins-Temporal Shift Module (BTSM) that shifts features along bins and temporal dimensions to fuse sparse events efficiently. We further establish the EgoEvGesture dataset, the first large-scale dataset for egocentric gesture recognition using event cameras. Experimental results demonstrate that our method achieves 62.7% accuracy tested on unseen subjects with only 7M parameters, 3.1% higher than state-of-the-art approaches. Notable misclassifications in freestyle motions stem from high inter-personal variability and unseen test patterns differing from training data. Moreover, our approach achieved a remarkable accuracy of 97.0% on the DVS128 Gesture, demonstrating the effectiveness and generalization capability of our method on public datasets. The dataset and models are made available at https://github.com/3190105222/EgoEv_Gesture.
@@ II-B Event-based Methods for Human Perception
TABLE I: Dataset comparison with key attributes
@@
Attributes: Modality (R=RGB, E=Event), 2H: Two-Hand, Res: Resolution (pixels), Real: Real-World, 1P: 1st-Person
@@ V-D Ablation Study
V-E Quantitative Result on DVS128 Gesture
@@
Our leading recognition performance on the DVS128 Gesture dataset stems from three key architectural innovations: (1) A lightweight CNN with asymmetric depthwise convolutions that maintains spatiotemporal features while reducing parameters, crucial for sparse event-based data; (2) A plug-and-play state-space context block that focuses on gesture dynamics; (3) The parameter-free BTSM module that effectively fuses sparse spatiotemporal events. Achieving 97.0% accuracy (matching S-tllr’s reimplemented performance) validates our architecture’s generalization capability and robustness for event-based gesture recognition.